# F-Fidelity: A Robust Framework for Faithfulness Evaluation of Explainable AI

### v1

@@ D.3 Fidelity and Explanation Size
D.4 ViT as Backbone for Image Classification.
To further validate the robustness of our F-Fidelity framework across different model architectures, we conducted additional experiments using ViT (Dosovitskiy et al., 2020) as the backbone for image classification on both Cifar-100 and TinyImageNet datasets. We utilize the SG-SQ (Smilkov et al., 2017) to generate explanations. For the settings, se follow our main experimental setup. As shown in Table 6, F-Fidelity consistently achieves best correlations across all metrics for both datasets, outperforming other methods. This strong performance on ViT models further emphasizes the versatility and effectiveness of F-Fidelity in evaluating explanations across different deep learning architectures.
Correlation Cifar-100 Tiny-Imagenet
D.5 Natural Language Classification Explanation Evaluation
Setup. We use two benchmark datasets for our NLP experiments: the Stanford Sentiment Treebank (SST) (Socher et al., 2013) for binary sentiment classification and the Boolean Questions (BoolQ) (Socher et al., 2013) dataset for question-answering tasks. For SST, we utilize 67,349 sentences for training and 872 for testing. BoolQ comprises 9,427 question-answer pairs for training and 3,270 for testing. We employ two popular model architectures: LSTM networks and Transformer-based models. For SST, the token embedding dimension is set to 128, and the hidden dimension is 512. The model has 2 multi-head attention layers, each with 4 heads. For BoolQ, the token embedding dimension is increased to 256, with 8 layers, each containing 8 heads. We follow the setting in image classification. Specifically, to generate explanations, we first use IG to obtain initial explanations, then apply noise perturbations to the importance of each timestamp at levels of $[0.0,0.2,0.4,0.6,0.8,1.0]$. We compare our F-Fidelity method against baselines including Fidelity (Fidelity) and R-Fidelity, evaluating performance using both macro and micro Spearman correlations. We use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of 1E-4, keeping other hyperparameters at their default values. The model accuracy results are shown in Table 7.
Table 7: The accuracy of the models to be explained in the natural language classification task.
LSTM Transformer
Dataset Origin Fine-tune Origin Fine-tune
SST2 0.8131 0.8280 0.7362 0.7867
BoolQ 0.6281 0.6263 0.6737 0.6764
Table 8: Spearman rank correlation and rank results on SST2 dataset with LSTM and Transformer.
Correlation LSTM Transformer
Fidelity R-Fidelity F-Fidelity Fidelity R-Fidelity F-Fidelity
Macro Corr. MoRF vs GT $\downarrow$ -1.$0$1.$0$1.$0$$0$1.$0$1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -1.$0$1.$0$1.$0$$0$1.$0$1.00$\pm$0.00
Micro Corr. MoRF vs GT $\downarrow$ -1.00$\pm$0.01 -1.00$\pm$0.01 -0.99$\pm$0.03 -0.93$\pm$0.07 -0.97$\pm$0.05 -1.00$\pm$0.01
LeRF vs GT $\uparrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.01 0.98$\pm$0.04 0.99$\pm$0.02 0.99$\pm$0.03
MoRF vs LeRF $\downarrow$ -1.00$\pm$0.01 -1.00$\pm$0.01 -0.99$\pm$0.02 -0.92$\pm$0.09 -0.97$\pm$0.05 -0.99$\pm$0.02
Micro Rank MoRF vs GT $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.05$\pm$0.22 2.05$\pm$0.89 1.32$\pm$0.46 1.05$\pm$0.22
LeRF vs GT $\downarrow$ 1.05$\pm$0.22 1.05$\pm$0.22 1.05$\pm$0.22 1.16$\pm$0.49 1.26$\pm$0.55 1.16$\pm$0.36
MoRF vs LeRF $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.11$\pm$0.31 2.05$\pm$0.97 1.32$\pm$0.46 1.11$\pm$0.45
Table 9: Spearman rank correlation and rank results on BoolQ dataset with LSTM and Transformer.
Correlation LSTM Transformer
Fidelity R-Fidelity F-Fidelity Fidelity R-Fidelity F-Fidelity
Macro Corr. MoRF vs GT $\downarrow$ -1.$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -1.$0$1.$0$1.$0$1.$0$1.$0$1.00$\pm$0.00
Micro Corr. MoRF vs GT $\downarrow$ -1.00$\pm$0.01 -1.00$\pm$0.01 -1.$0$1.00$\pm$0.01 -1.00$\pm$0.01 -1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ -1.00$\pm$0.01 -1.00$\pm$0.01 -1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
Micro Rank MoRF vs GT $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
LeRF vs GT $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
MoRF vs LeRF $\downarrow$ 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
Results. Our experiments in the NLP domain reveal an interesting phenomenon that NLP models demonstrate remarkable robustness to OOD inputs, which is reflected in the performance of various fidelity metrics. Notably, the vanilla Fidelity (Fidelity) method achieves excellent results, often matching or closely approaching the performance of more sophisticated methods like R-Fidelity and our proposed F-Fidelity. This is evident from the near-perfect correlations (-1.00 or 1.00) across most metrics for both LSTM and Transformer models. This strong performance of vanilla fidelity suggests that the OOD problem, which often necessitates more complex evaluation frameworks in other domains, may be less pronounced in NLP tasks. The robustness of NLP models to perturbations in input tokens likely contributes to this phenomenon, allowing simpler evaluation methods to maintain their efficacy. However, it’s worth noting that our proposed F-Fidelity method still demonstrates consistent top-tier performance across all metrics and models, reinforcing its versatility and reliability even in scenarios where simpler methods perform well.
@@ E.1 Image Classification Explanation Evaluation
Experiments on Tiny-Imagenet. Similar to the Cifar-100, we provide the detailed experiment results on Tiny-Imagenet in this section, which is a scaled-down version of the ImageNet dataset, designed for image classification tasks. Tiny-Imagenet contains 200 classes and 500 training images per class, 50 validation images, and 50 test images per class. The image resolution is 64x64. From Table 12 and  13, we observe that compared to R-Fidelity and F-Fidelity, F-Fidelity and R-Fidelity have the best performance.
@@ E.2 Detailed Results on Time series Datasets
We conduct the experiments on the PAM and Boiler datasets with LSTM architecture. As Table 14 and  15 show, our method distinguishes explanations with a larger margin than R-Fidelity. Moreover, Fidelity and ROAR method fail to distinguish different explanations in the Boiler dataset.
@@ E.3 Detailed Results on NLP
In Table 16,  1718 and  19, with LSTM and Transformer, we find the Fidelity has a consistent result which is quite different than other tasks. Compared to R-Fidelity and F-Fidelity, the results are almost the same both in the SST2 and BoolQ datasets, which indicates in these two cases, the OOD problem can be ignored.
@@ E.4 Detailed results of hyperparameter sensitivity study
In this section, we provide the detailed ablation study experiment results on the Colored-Mnist dataset. The experiment settings are described as Section D.2. We provide the detailed $Fid^{+}$, $Fid^{-}$, and three Spearman rank Correlation, same as the subsection E.1. As Table 2021, and  22 show, our method is robust to the choice of hyperparameters. From the $Fid^{+}$ and $Fid^{-}$, the larger sampling number helps reduce the standard deviation. But with large sampling number, the evaluation time increases. In this paper, we select a balance point of $50$ for all the other experiments. For the hyperparameter $\beta$, the best option is near the ground truth explanation size ratio. However, in the real-world setting, it is hard to find the best choice. In our paper, for the fine-tuning efficiency and evaluation performance, we select $0.1$ for all the other experiments. In R-Fidelity, the hyperparameter $\alpha$ is designed for alleviating the OOD problem, where smaller $\alpha$ means less OOD problem. In our method, because of the $\beta$, a larger $\alpha$ is also optional.
Table 20: The detailed ablation study results on sampling number $T$, with $\beta=0.1$, $\alpha_{1}=\alpha_{2}=0.5$.
Sampling Number $T$ $Fid^{+}$ $Fid^{-}$ Local Spearman Correlation
@@
Table 21: The detailed ablation study results on hyperparameter $\beta$, with $S=50$, $\alpha_{1}=\alpha_{2}=0.5$.
@@
Table 22: The detailed ablation study results on hyperparameter $\alpha=\alpha^{+}=\alpha^{-}$, with $T=50$, $\beta=0.1$.

### v2

@@ D.8 Emprical Verification of determining the Explanation Size
D.9 Hyperparameter Sensitivity Study
In this part, we assess the robustness of F-Fidelity to hyperparameter choices. We select a subset of 400 samples with an explanation size ratio of 0.2 from the colored-MNIST dataset (Arjovsky et al., 2019). we have three key hyperparameters: the number of sampling iterations $N$, the ratio $\alpha^{+}/\alpha^{-}$, and $\beta$. For each hyperparameter, we explored a range of values to understand their impact on the performance of F-Fidelity. We evaluated the method’s performance using three Macro Spearman correlations, with additional detailed results on $RFid^{+}$, $RFid^{-}$, and Micro Spearman correlations provided in Appendix E.4.
(a) Ablation study on $N$
(b) Ablation study on $\beta$
(c) Ablation study on $\alpha$
Figure 5: Ablation study on sampling number $N$, $\beta$, and $\alpha=\alpha^{+}=\alpha^{-}$. We report the Macro Spearman rank correlations.
D.10 Comparing Different Explanation Methods
To demonstrate F-Fidelity’s ability to differentiate between explanation methods of varying quality, we conduct a comparative study between GradCAM and Saliency Map (Simonyan et al., 2013) on both Tiny-Imagenet and ImageNet. GradCAM typically produces more focused explanations by utilizing class-specific gradient information at higher-level feature maps, while Saliency Map operates directly on pixel-level gradients.
Our evaluation on Tiny-Imagenet confirms this expected difference. Across our test set, GradCAM achieves better performance with higher $FFid^{+}$ (0.2286 vs 0.2020) and lower $FFid^{-}$ (0.0547 vs 0.1050) scores than Saliency Map. Note that both explanation methods are used to generate explanations with respect to the ground truth label. Thus, the FFid values are negative if the original prediction is incorrect.
Table 11: Comparative evaluation of GradCAM and Saliency Map explanations using F-Fidelity on Tiny-Imagenet. We show five cases with their original images and corresponding explanation visualizations. The $FFid^{+}$ and $FFid^{-}$ scores are averaged over different sparsity levels, with bold scores indicating better performance. The $FFid^{+}$ and $FFid^{-}$ scores of case 5 are negative because the original image is misclassified. The bottom row reports metrics averaged across all samples in the test set.
Image Explanation Visualization Evaluation Metrics
Original GradCAM Saliency Map Method $FFid^{+}\uparrow$ $FFid^{-}\downarrow$
Case 1 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.8743
0.8638
0.1267
0.3342
Case 2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.3838
0.3143
0.029
0.0048
Case 3 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.2095
0.0752
0.0190
0.0314
Case 4 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.4552
0.1095
0.0486
0.1000
Case 5 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
-0.0057
-0.0086
-0.0048
-0.0029
Average over All Images in The Test Set
GradCAM
Saliency Map
0.2286
0.2020
0.0547
0.1050
To verify our method’s effectiveness on high-dimensional data, we further evaluate on a subset of ImageNet (5000 images, 224×224×3 resolution) using ResNet-18 as the backbone. Unlike our previous experiments where we fine-tuned on training data and evaluated explainers on a separate test set, here we fine-tune and evaluate on the same image set to enable controlled comparison between explainers. Note that this fine-tuned model is specifically optimized for this dataset and should only be used to compare explainer performance within this set. Following the same protocol as our Tiny-Imagenet experiments, we fine-tune the model using Adam optimizer with default settings and evaluate on 500 sampled images from this set. As shown in Table 12, The results consistently show GradCAM outperforming Saliency Map, with higher $FFid^{+}$ (0.1123 vs 0.0672) and lower $FFid^{-}$ (0.0046 vs 0.0091) scores. It shows that F-Fidelity successfully captures the qualitative differences between explainers, such as GradCAM’s more focused attribution maps compared to Saliency Map’s noisier explanations.
Table 12: Comparative evaluation of GradCAM and Saliency Map explanations using F-Fidelity on ImageNet dataset with ResNet-18. The $FFid^{+}$ and $FFid^{-}$ scores are averaged over different sparsity levels, with bold scores indicating better performance.
Image Explanation Visualization Evaluation Metrics
Original GradCAM Saliency Map Method $FFid^{+}\uparrow$ $FFid^{-}\downarrow$
Case 1 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.3476
0.2362
0.0019
0.0057
Case 2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.3981
0.3571
0.0076
0.0105
Case 3 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.5438
0.2162
0.0257
0.0590
Case 4 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
0.6190
0.1019
0.0029
0.0162
Case 5 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GradCAM
Saliency Map
-0.0010
-0.1477
-0.2467
-0.0895
Average Over 500 Images
GradCAM
Saliency Map
0.1123
0.0672
0.0046
0.0091
D.11 Computational Efficiency Analysis
Following the comparative setup between GradCAM and Saliency Map on Tiny-Imagenet, we conduct a detailed computational efficiency analysis. Table 13 breaks down the computational time requirements for both ROAR and F-Fidelity. The key advantage of F-Fidelity lies in its one-time fine-tuning cost - while ROAR requires separate explaining and training processes for each explainer (totaling 7140s for two explainers), F-Fidelity needs only a single fine-tuning process (2925s) that can be reused across all explainers.
For the Tiny-Imagenet dataset with ResNet backbone, ROAR’s per-explainer cost includes generating explanations for training samples (approximately 650-700s), model retraining (around 2900s), and evaluation (about 120s). In contrast, F-Fidelity requires only a single fine-tuning step (2925s), followed by evaluation for each explainer (approximately 550s). This translates to a total processing time of 4026s for F-Fidelity compared to 7378s for ROAR when evaluating two explainers, with the gap widening further as more explainers are evaluated.
This efficiency gain becomes particularly significant when evaluating multiple explainers or conducting extensive ablation studies, making F-Fidelity more practical for real-world applications while maintaining superior evaluation quality as demonstrated in our previous experiments. The advantage is even more pronounced for computationally intensive explainers, as ROAR requires generating explanations for the entire training set (100k+ samples in our case), while F-Fidelity only needs explanations for the evaluation set. Moreover, our hyperparameter analysis shows that we can further reduce the evaluation time by decreasing the number of samples - with $N=10$, the evaluation time per explainer drops from approximately 550s to 180s while maintaining robust performance. This flexibility allows users to balance between evaluation speed and precision based on their specific needs.
Table 13: Computational time analysis for evaluating GradCAM and Saliency Map on Tiny-Imagenet using ResNet as the backbone. While ROAR requires separate explaining and training processes for each explainer, F-Fidelity’s fine-tuning cost is one-time.
Method Explaining Training Model Training/ Evaluation Total
Samples (100k) Fine-tuning (100 epochs) (2000 samples) Time
ROAR 647s + 721s 2911s + 2861s 114s+124s 7378s
F-Fidelity - 2925s 541s+560s 4026s
@@ E.1 Image Classification Explanation Evaluation
Experiments on Tiny-Imagenet. Similar to the CIFAR-100, we provide the detailed experiment results on Tiny-Imagenet in this section, which is a scaled-down version of the ImageNet dataset, designed for image classification tasks. Tiny-Imagenet contains 200 classes and 500 training images per class, 50 validation images, and 50 test images per class. The image resolution is 64x64. From Table 16 and  17, we observe that compared to R-Fidelity and F-Fidelity, F-Fidelity and R-Fidelity have the best performance.
@@ E.2 Detailed Results on Time series Datasets
We conduct the experiments on the PAM and Boiler datasets with LSTM architecture. As Table 18 and  19 show, our method distinguishes explanations with a larger margin than R-Fidelity. Moreover, Fidelity and ROAR method fail to distinguish different explanations in the Boiler dataset.
@@ E.3 Detailed Results on NLP
In Table 20,  2122 and  23, with LSTM and Transformer, we find the Fidelity has a consistent result which is quite different than other tasks. Compared to R-Fidelity and F-Fidelity, the results are almost the same both in the SST2 and BoolQ datasets, which indicates in these two cases, the OOD problem can be ignored.
@@ E.4 Detailed results of hyperparameter sensitivity study
In this section, we provide the detailed ablation study experiment results on the Colored-Mnist dataset. The experiment settings are described as Section D.9. We provide the detailed $Fid^{+}$, $Fid^{-}$, and three Spearman rank Correlation, same as the subsection E.1. As Table 2425, and  26 show, our method is robust to the choice of hyperparameters.
Table 24: The detailed ablation study results on sampling number $N$, with $\beta=0.1$, $\alpha^{+}=\alpha^{-}=0.5$.
Sampling Number $N$ $Fid^{+}$ $Fid^{-}$ Local Spearman Correlation
@@
Table 25: The detailed ablation study results on hyperparameter $\beta$, with $N=50$, $\alpha^{+}=\alpha^{-}=0.5$.
@@
Table 26: The detailed ablation study results on hyperparameter $\alpha=\alpha^{+}=\alpha^{-}$, with $N=50$, $\beta=0.1$.